OpenStudy (anonymous):

Solve for x. 5^(2x)-5^(x+1)+6=0

OpenStudy (anonymous):

i got x=log base 5 of 3 and x=log base 5 of 2....are both answers right?

OpenStudy (anonymous):

Use a variable Y=5^x and solve for Y, then solve for x like x=log(Y)/log(5)

OpenStudy (anonymous):

i used the variable but then i factored to get (y-3) (y-2)=0 and got the answer above. but i dont have a calculator with me to check both answers

OpenStudy (anonymous):

leave it like x1=log(3)/log(5) and x2=log(2)/log(5) for your solutions for Y are 3 and 2

OpenStudy (anonymous):

so the answers are 3 and 2? sorry that kinda confused me.

OpenStudy (anonymous):

those are the solutions for Y, now you have to calculate x. As you know log(Y)=log(5^x)=x*log(5), then x=log(Y)/log(5) note:logarithms can be calculated in any base, say 10, neperian...

OpenStudy (anonymous):

oh ok i think that makes some more sense haha. thank you so much for your help

OpenStudy (anonymous):

u r welcome